### v1

@@ 4 Experiments
Our proposed framework is primarily focused on advancing UQ in time-series exhibiting volatility clustering. In this section, we detail experiment results in our motivating application — time-series forecasting and UQ on cryptocurrency and U.S. equities time-series datasets, to illustrate the benefits of our proposed method. Nonetheless, SMD parameterization, modelling distribution parameters using subnetworks and ensemble predictions can also be applied to general applications of prediction UQ. In Appendix A.1, we also compare our method to Ensemble and Evidential using the UCI benchmark dataset. This is intended to provide readers with a direct comparison to the results published in Lakshminarayanan et al., (2017) and Amini et al., (2020), demonstrating the benefits of our proposed improvements in non-time-series datasets.
4.1 Uncertainty quantification in financial time-series forecasting
@@
Our U.S. equities experiment follows the same setup as cryptocurrencies. Mimicking the S&P 500 index universe, the dataset consists of daily returns downloaded from the Wharton Research Data Service over 1984 to 2020, for the 500 largest stocks [^9] The list of stocks is refreshed every June, keeping the same stocks until the next rebalance. listed on NASDAQ, NYSE and NYSE American. Data from 1984 to 1993 are used for hyperparameter tuning, while 1994 to 2020 are used for out-of-sample testing. The network is refitted every January using a rolling 10-year window. Each input sequence consists of 240 trading days (approximately one-year) of daily returns $r$ and squared returns $\log(r^{2})$, forecasting forward 20-day (approximately one-month) return and its uncertainty. Note that the unit of analysis is monthly and unit of observation is daily. One-month is a popular forecast horizon for U.S. equities in literature (e.g., Gu et al.,, 2020 and Wong et al.,, 2021), which motivated our choice of forecast horizon.
At this point, it is useful to remind readers that prior literature have found both datasets to exhibit time-varying variance (e.g., Cont,, 2001; Hafner,, 2018), which is also visible in Figure 2. We start with the main empirical results of this work. Table 2 contains forecast results on both the cryptocurrency dataset (left) and U.S. equities (right). We observe that Combined has the highest average cross-sectional correlation (higher is better), and lowest RMSE and NLL (both lower is better) in both datasets. This indicates that Combined has higher cross-sectional predictive efficacy (as measured by correlation) and is able to better forecast uncertainty of the time-series prediction. Evidential has higher (better) correlation and lower RMSE (better) than Ensemble but higher (worse) NLL in cryptocurrency. In U.S. equities, Evidential has worse correlation, RMSE and NLL than Ensemble. Correlations in U.S. equities are materially lower for all three methods compared to the cryptocurrency dataset. We hypothesize that this is due to both the difference in forecast horizon and maturity of the U.S. stock market.
Table 2: Main results: Comparing Ensemble, Evidential and Combined on average cross-sectional correlation, RMSE and NLL for cryptocurrencies (left) and U.S. equities (right), respectively. Average result and standard deviation over 10 trials for each method. Best method for each dataset is highlighted in bold.
@@
Figure 2: Forecast uncertainties of Combined, Ensemble and Evidential applied to Bitcoin (left) and Chevron Corp (right), respectively. For Bitcoin, square root of the average over each day shown.
@@
Next, Figure 2 compares predicted uncertainty and actual prediction error of the three methods to actual volatility of Bitcoin (BTC/USDT), the cryptocurrency with the highest market capitalization, and Chevron Corp., a major U.S. oil producer which have endured multiple market shocks. Volatility forecasts are often compared with observed volatility (typically computed over a look back window) to evaluate forecast performance. However, the true instantaneous volatility of an asset (i.e., $\sigma^{2}$ in Equation (1)) is unobservable (Ge et al.,, 2022). Thus, in the top row of Figure 2, we use the standard deviation of hourly returns computed over each day for Bitcoin and absolute value of monthly returns for Chevron (as forecast horizon is monthly), as proxies for $\sigma^{2}$. In rows 2–4, for Bitcoin, we aggregate hourly forecasts to daily data points by computing the daily root mean squared return forecast error $\sqrt{\tfrac{1}{24}\sum_{k=0}^{23}(y_{t-k}-\hat{y}_{t-k})^{2}}$ (denoted $\sqrt{(y-\hat{y})^{2}}$) computed from hourly return forecasts, and the daily root mean predictive uncertainty $\sqrt{\tfrac{1}{24}\sum_{k=0}^{23}\operatorname{Var}(\hat{y}_{t-k})}$ (denoted $\sqrt{\operatorname{Var}(\hat{y})}$), for each $t=24,48,72,\dots,T$ (note that $t$ for cryptocurrency is in hourly units). For Chevron, as the unit of analysis is monthly, we plot the absolute error between actual monthly returns and predicted returns (denoted $|y-\hat{y}|$), and square-root of forecast uncertainty (denoted $\sqrt{\operatorname{Var}(\hat{y})}$) on the right column of Figure 2. Comparing the top row of Figure 2 to the root return forecast error of row 2-4 (blue line), we observe that forecast error spikes when volatility of the asset spikes. This is expected, as the spike in volatility leads to large forecast errors. Comparing the bottom three rows of Figure 2, which correspond to Combined, Ensemble and Evidential, respectively. We observe that Combined’s predicted uncertainty of $\hat{\mu}$ tracks actual forecast error much more closely than Evidential and Ensemble. This appears to be especially true during periods of elevated volatility, which are important to investors. Overestimation of predictive uncertainty is severe for Ensemble in Bitcoin, where predictive uncertainty can sometimes be significantly higher than observed forecast error. Lastly, Evidential underestimates forecast error during heightened volatility (e.g., March 2020 in Figure 2) and overestimates forecast error under periods of low volatility (e.g., July 2020 in Figure 2).
@@
For Chevron, we observe multiple spikes of volatility, corresponding to the U.S. recession over 2008-09, the oil shock in 2015 and the 2020 pandemic. Notably, we observe that Combined tracks the spike in forecast error better than Ensemble and Evidential. Note that the “block-like” appearances of uncertainty forecasts of both Ensemble and Evidential are due to periodic training (monthly for cryptocurrencies and yearly for U.S. equities) and the failure to generalize the prevailing volatility environment. During training, the optimizer updates network weights $\bm{W}$ and bias $\bm{b}$ (which is analogous to the intercept in linear models). When the network fails to generalize, it minimizes the loss function by updating the bias rather than the weights. Thus, outputting the same constant that do not vary with the input, until the network is re-trained in the following month (for cryptocurrencies) or year (for U.S. equities). This produces the block-like appearances of Ensemble and Evidential, and is indicative of the network setup (e.g., no separate modelling of hyperparameters) being unsuitable to this class of problems. We observe similar visual characteristics in the predicted uncertainty of other cryptocurrencies and stocks.
@@ 4.2 Ablation study
Table 3: Ablation studies of Combined: In each column, we remove model averaging (No Averaging), separate modelling of distribution parameters (Single Output) and using return time-series only (Returns-only) for cryptocurrencies (left) and U.S. equities (right), respectively. Average result and standard deviation over 10 trials for each method. Note that cryptocurrency returns are hourly and U.S. stock returns are monthly.
@@
[…]
@@ 5 Conclusions
Our motivating application of portfolio selection depends on both forecasts and forecast uncertainties. This is a challenging problem due to both the low signal-to-noise ratio in financial markets (Gu et al.,, 2020) and the presence of volatility clustering. To this end, we present a method for the simultaneous forecasting asset returns and modelling of forecast uncertainty in presence of volatility clustering. Our proposed method extends and simplifies the work of Lakshminarayanan et al., (2017) and Amini et al., (2020). We propose to use a SMD (which uses a Gamma prior for scale uncertainty $\nu$) as a simpler alternative to a NIG prior, in which a Normal prior is placed on $\mu$ and an Inverse-Gamma prior on $\sigma^{2}$). Parameters of SMD are modelled using separate subnetworks. Together with ensembling and the use of second order of returns as inputs, we show that our proposed method can successfully model time-varying variance of the DGP, while providing superior forecasting performance than two state-of-the-art neural network UQ methods — Evidential and Ensemble. This is illustrated through the successful quantification of forecast uncertainty of two financial time-series datasets: cryptocurrency and U.S. equities.
@@
Our proposed SMD formulation offers an avenue to resolve some of the criticisms of Meinert et al., (2022) and Bengs et al., (2023). In particular, our SMD parameterization has three effective parameters and thus does not have any unresolved degrees of freedom. We can set $\alpha=\beta$, which leads to a marginal t-distribution where the three distributional parameters ($\gamma,\sigma^{2},\alpha$) relate directly to the location, scale and shape of the t-distribution, without the need of a hierarchical model. In this formulation, epistemic uncertainty is assumed to be the difference between the predictive (t-distributed) and aleatoric (Normal-distributed) uncertainties. This assumption provides for a simpler model but lacks the granular attribution between aleatoric and epistemic uncertainties afforded by the NIG prior in Evidential. However, as Meinert et al., (2022) has pointed out, the granular control comes at the cost of an unresolved degree of freedom. Thus, users are encouraged to weigh the trade-offs in choosing a method to deploy. This also makes for a potential future research direction. We show empirically that our method is able to accurately predict forecast errors, similar to the success Evidential demonstrated in other real world applications (e.g., see Liu et al.,, 2021; Soleimany et al.,, 2021; Cai et al.,, 2021; Singh et al.,, 2022; Li and Liu,, 2022). From a finance application perspective, forecast uncertainty can be used to size bets, or as advanced warning to protect the portfolio from downside risk. For example, if forecast uncertainty reaches a certain threshold, an investor could purchase portfolio insurance (e.g., put options) or liquidate positions to reduce risk. The ability to attribute epistemic and aleatoric uncertainties may also allow for more advanced portfolio optimization techniques to be developed in future research (e.g., place different risk aversions on the two sources of uncertainties). Lastly, UQ in time-series applications is a relatively under-explored area of literature. We believe this work can lead to further advancements of UQ in complex time-series.
@@ Appendix A Marginal distribution of a Scale Mixture
From Equation (10), we have $\operatorname{N}(y|\gamma,\tfrac{\sigma^{2}}{\lambda})\operatorname{Gam}(\lambda|\alpha,\beta)$. Marginalizing over $\lambda$ produces the data likelihood,
$\mathrm{p}(y|\gamma,\sigma^{2},\alpha,\beta)$ $=\int_{\lambda}^{\infty}\mathrm{p}_{\operatorname{N}}(y|\gamma,\sigma^{2}\lambda^{-1})\mathrm{p}_{\operatorname{G}}(\lambda|\alpha,\beta)\mathop{}\!\mathrm{d}\lambda$
$=\int_{\lambda}^{\infty}\left[\sqrt{\frac{\lambda}{2\pi\sigma^{2}}}\exp\left\{-\frac{\lambda(y-\gamma)^{2}}{2\sigma^{2}}\right\}\right]\left[\frac{\beta^{\alpha}}{\Gamma(\alpha)}\lambda^{\alpha-1}\exp^{-\beta\lambda}\right]\mathop{}\!\mathrm{d}\lambda$
$=\frac{\beta^{\alpha}}{\Gamma(\alpha)\sqrt{2\pi\sigma^{2}}}\int_{\lambda=0}^{\infty}\lambda^{\alpha-\tfrac{1}{2}}\exp\left\{-\frac{\lambda(y-\gamma)^{2}}{2\sigma^{2}}-\beta\lambda\right\}\mathop{}\!\mathrm{d}\lambda$
$=\frac{\beta^{\alpha}}{\Gamma(\alpha)\sqrt{2\pi\sigma^{2}}}\left[\frac{(y-\gamma)^{2}}{2\sigma^{2}}+\beta\right]^{-(\alpha+\tfrac{1}{2})}\int_{\lambda=0}^{\infty}\left\{\lambda\left[\frac{(y-\gamma)^{2}}{2\sigma^{2}}+\beta\right]\right\}^{\alpha-\tfrac{1}{2}}$
$\quad\exp\left\{-\lambda\left[\frac{(y-\gamma)^{2}}{2\sigma^{2}}+\beta\right]\right\}\mathop{}\!\mathrm{d}\left\{\lambda\left[\frac{(y-\gamma)^{2}}{2\sigma^{2}}+\beta\right]\right\},$
@@
$=\frac{\Gamma(\alpha+\tfrac{1}{2})}{\Gamma(\alpha)}\frac{1}{\sqrt{2\pi\sigma^{2}\beta}}\left[\frac{(y-\gamma)^{2}}{2\sigma^{2}\beta}+1\right]^{-(\alpha+\tfrac{1}{2})}\hskip 119.50157pt$
@@
A.1 Benchmarking on UCI dataset
Table 4: Comparing Ensemble (Lakshminarayanan et al.,, 2017), Evidential (Amini et al.,, 2020) and Combined (this work) on RMSE and NLL using the UCI benchmark datasets. Average result and standard deviation over 5 trials for each method. The best method for each dataset and metric are highlighted in bold.

### v2

@@ 3 Experiments
Our proposed framework is primarily focused on advancing UQ in time-series exhibiting volatility clustering. In this section, we detail experiment results in our motivating application — time-series forecasting and UQ on cryptocurrency and U.S. equities time-series datasets, to illustrate the benefits of our proposed method. Nonetheless, SMD parameterization, subnetwork construction for each distribution parameter and ensemble predictions can also be applied to general applications of predictive UQ. In Appendix B, we also compare our method to Ensemble and Evidential methods using the UCI benchmark dataset. This is intended to provide readers with a direct comparison to the results published in Lakshminarayanan et al., (2017) and Amini et al., (2020), demonstrating the benefits of our proposed combined method in non-time-series datasets.
3.1 Applications in financial time-series
In this section, we will first apply the three methods on the cryptocurrency dataset, then the U.S. equities dataset [^11] Note that in this section, “forecast uncertainty” and “uncertainty forecast” refer to estimated predictive uncertainty (i.e., sum of epistemic and aleatoric uncertainties) for simplicity.. Prior literature have found both datasets to exhibit time-varying variance (e.g., Cont,, 2001; Hafner,, 2018) and so they are suitable for demonstrating our proposed Combined method in comparison with Ensemble and Evidential methods. To compare performance, we use mean cross-sectional correlation (CC) and root mean square error (RMSE),
@@ 3.1 Applications in financial time-series
where $\bm{y}_{t}$ and $\hat{\bm{y}_{t}}$ are vectors of target and prediction at time $t$ during the test period and $t$ is hourly for cryptocurrencies and monthly for U.S. equities as measures of predictive accuracy, in addition to NLL as a measure of model fit (Ambachtsheer,, 1974; Grinold and Kahn,, 1999; Wong et al.,, 2021). The same neural network architectures are used in the two datasets, with hyperparameters tuned independently.
3.1.1 Cryptocurrency dataset
[…]
A network is trained every 30 days using an expanding window of all cryptocurrency data from July 2018. Each input sequence consists of 10 days of hourly returns $r$ and log squared returns $\log(r^{2})$ (i.e., the input is a matrix with dimensions $240\times 2$), and are used to predict forward one hour return (i.e., units of analysis and observation are both hourly). This training scheme is shared across all three models in Table 1. Network topology consists of LSTM layers, followed by fully connected layers with ReLU activation and the respective output layers of Ensemble (Gaussian) and Evidential (NormalInverseGamma). For Combined, we use four separate subnetworks as illustrated in Figure 2. As discussed in Section 1.4, we consider UQ in cryptocurrencies to be especially challenging due to their high volatility.
3.1.2 U.S. equities dataset
[…]
3.1.3 Empirical results
Table 2 reports forecast results for both cryptocurrency (left) and U.S. equities (right) datasets. We observe that Combined method has the highest average cross-sectional correlation (higher is better), and lowest RMSE and NLL (both lower is better) in both datasets. This indicates that Combined method has higher cross-sectional predictive efficacy (as measured by correlation) and is able to better forecast uncertainty of the time-series prediction. In cryptocurrency, Evidential method has higher (better) correlation, lower RMSE (better) but higher (worse) NLL than Ensemble method. In U.S. equities, Evidential method has worse correlation, RMSE and NLL than Ensemble method. Correlations in U.S. equities are materially lower for all three methods compared to the cryptocurrency dataset. We hypothesize that this is due to both the difference in forecast horizon and maturity of the U.S. stock market.
Table 2: Main results: Comparing Ensemble, Evidential and Combined methods on average cross-sectional correlation, RMSE and NLL for cryptocurrencies (left) and U.S. equities (right), respectively. Average result and standard deviation are computed over 10 trials for each method. Best method for each dataset is highlighted in bold.
@@ 3.1.3 Empirical results
Figure 3: Forecast uncertainties of Combined, Ensemble and Evidential applied to Bitcoin (left; daily) and Chevron Corporation (right; monthly), respectively over the out-of-sample period. For the top row, SD for Bitcoin and AV for Chevron in Equation (21) are shown.
To compare relative UQ performance, we apply the trained networks to Bitcoin (BTC/USDT), the cryptocurrency with the highest market capitalization, and Chevron Corporation, a major U.S. oil producer which have endured multiple market shocks. Additionally, we use the standard deviation (SD) of hourly returns computed over each day for Bitcoin, and absolute value (AV) of monthly returns $|y_{t}|$ for Chevron to provide visual confirmation of volatility clustering, given respectively by,
$\text{SD}=\sqrt{\frac{1}{24}\sum_{k=0}^{23}(y_{t-k}-\bar{y}_{t-k})^{2}}\quad\text{and}\quad\text{AV}=|y_{t}|,$ (21)
where $t$ is hourly and monthly time. Figure 3 visualizes these volatility clustering in the top row. As expected, Bitcoin displays sharper spikes of clustered volatility due to its more volatile nature. For Chevron, we observe multiple spikes of volatility, corresponding to the U.S. recession over 2008-09, the oil shock in 2015 and the 2020 pandemic. These spikes are relatively lower than those of Bitcoin due to the lower frequency of return forecast for Chevron (monthly) relative to Bitcoin (hourly). Hence, the dynamics of returns volatility differ across horizon of returns.
Next, we evaluate the performance of predictive uncertainty forecast during the test period. Predictive uncertainty are given by $\operatorname{Var}[\hat{y}]=\hat{\sigma^{2}}$ for Ensemble in Equation (3), $\operatorname{Var}[y]=\tfrac{\hat{\beta}(1+\hat{\nu})}{\hat{\nu}(\hat{\alpha}-1)}$ for Evidential method in (10) and $\operatorname{Var}[\hat{y}]=\operatorname{Var}[\tilde{y}]$ for Combined method in (20). They should then be compared with observed volatility over a look back window. However, the true instantaneous volatility of an asset (i.e., $\sigma^{2}$ in Equation (1)) is not directly observable (Ge et al.,, 2022). Hence, we estimate the true volatility using different formulations of actual forecast error for the returns of Bitcoin and Chevron Corporation as they have different units of analysis. For Bitcoin, we compute the daily root mean squared return forecast error (RMSE) using hourly return forecast as the true daily volatility proxy and compare it to daily root mean predictive uncertainty (RMPU) given by respectively,
$\text{RMSE}=\sqrt{\frac{1}{24}\sum_{k=0}^{23}(y_{t-k}-\hat{y}_{t-k})^{2}}\quad\text{and}\quad\text{RMPU}=\sqrt{\frac{1}{24}\sum_{k=0}^{23}\operatorname{Var}(\hat{y}_{t-k})},$
for each $t=24,48,72,\dots,T$ where $t$ for cryptocurrency is in hourly units. The two measures, true proxy and predictive uncertainty, are denoted by $\sqrt{(y-\hat{y})^{2}}$ and $\sqrt{\operatorname{Var}(\hat{y})}$ respectively. For Chevron Corporation, as the forecast horizon is monthly, we adopt the absolute monthly return forecast error (AE) between actual monthly returns and predicted returns (denoted by $|y-\hat{y}|$) as a proxy for the true $\sigma^{2}$ and the square-root of forecast uncertainty (MAPU; denoted by $\sqrt{\operatorname{Var}(\hat{y})}$). We note that SD for Bitcoin or AV for Chevron in Equation (21) are very similar to RMSE for Bitcoin and AE for Chevron since $\hat{y}\approx\bar{y}\approx 0$.
@@
Figure 3 compares these predictive uncertainty $\big{(}\sqrt{\operatorname{Var}(\hat{y})}\big{)}$ (orange line) with actual prediction errors ($\sqrt{(y-\hat{y})^{2}}$ or $|y-\hat{y}|$; blue line) for the three methods (row 2-4) applied to Bitcoin (left column) and Chevron Corporation (right column). We observe in these plots that predictive uncertainty spikes when the actual forecast error of the asset spikes. This is expected, as the spike in volatility (measured by actual forecast errors) leads to large predictive uncertainty. Comparing row 2-4 of Figure 3, which correspond to Combined, Ensemble and Evidential, respectively, we observe that Combined’s predicted uncertainty of $\hat{\mu}$ tracks actual forecast error much more closely than Evidential and Ensemble. This appears to be especially true during periods of elevated volatility, which are important to investors. Overestimation of predictive uncertainty is severe for Ensemble in Bitcoin, where predictive uncertainty can sometimes be significantly higher than actual forecast error. Lastly, Evidential underestimates forecast error during heightened volatility (e.g., March 2020 in Figure 3) and overestimates forecast error under periods of low volatility (e.g., July 2020 in Figure 3). On the other hand, we also observe that Combined method tracks the spike in forecast error better than Ensemble and Evidential.
@@
We note that the “block-like” appearances of uncertainty forecasts of both Ensemble and Evidential are due to periodic training (monthly for cryptocurrencies and yearly for U.S. equities) and the failure to generalize the prevailing volatility environment. During training, the optimizer updates network weights $\bm{W}$ and bias $\bm{b}$ (which is analogous to the intercept in linear models). When the network fails to generalize, it minimizes the loss function by updating the bias rather than the weights. Thus, outputting the same constant that do not vary with the input, until the network is re-trained in the following month (for cryptocurrencies) or year (for U.S. equities). This produces the block-like appearances of Ensemble and Evidential, and is indicative of the network setup (e.g., no separate modelling of hyperparameters) being unsuitable to this class of problems. We observe similar visual characteristics in the predicted uncertainty of other cryptocurrencies and stocks.
(a) Actual & Estimated Returns of Bitcoin
(b) Actual & Estimated Returns of Chevron
Figure 4: Actual returns and mean estimates of Combined, Ensemble and Evidential applied to Bitcoin (left; hourly) and Chevron (right; monthly), respectively. For legibility, we have shown hourly forecasts on March 14, 2020, and monthly forecasts over 2019-2020 for Bitcoin and Chevron, respectively.
Lastly, in Figure 4, we compare the mean estimates of Combined, Ensemble and Evidential to actual realized returns. To improve legibility, we have shown hourly forecasts for Bitcoin on March 14, 2020, a day of high volatility in cryptocurrencies, and monthly forecasts for Chevron over 2019-2020. The top row of Figure 4 compares mean estimates of the three models to actual realized returns of Bitcoin (left) and Chevron (right). Due to the low signal-to-noise ratio in financial data, the mean estimates have materially smaller scales than actual returns for all three models. Next, focusing on the bottom row of Figure 4, we observe higher variability for Combined in Bitcoin than Ensemble and Evidential. Evidential produced estimates that were persistently negative over this 24-hour window and did not anticipate the fluctuations of actual returns. Turning to Chevron, all three methods predict positive returns throughout 2019 but the fluctuations in estimates appear to track the dip in April and jump in June. However, all three methods could only forecast the sharp fall and subsequent reversal in returns over March and April 2020 at a lag of one month. This is to be expected as the volatility was due to an exogenous shock (pandemic).
@@ 3.2 Ablation study
[…]
Table 3: Ablation studies of Combined method: In each column, we remove model averaging (No Averaging), separate modelling of distribution parameters (Single Output) and using return time-series only (Returns-only) for cryptocurrencies (left) and U.S. equities (right), respectively. Average result and standard deviation over 10 trials for each method. Note that cryptocurrency returns are hourly and U.S. stock returns are monthly.
@@
Using a single output layer for all distribution parameters also leads to marginally worse NLL. Correlation is lower in cryptocurrencies but marginally higher in U.S. equities. While using returns only leads to marginally higher correlation but marginally worse on NLL in cryptocurrency, and lower correlation and worse NLL in U.S. equities. From Figure 5, the block-like appearances indicate that both using single output layer and using returns only result in the network failing to closely track time-varying variance of the DGP. This suggests that both squared returns and separate modelling of distribution parameters are required to model time-varying forecast uncertainty.
@@ 4 Conclusions
Our motivating application of portfolio selection depends on both return forecasts and uncertainties of return forecasts. This is a challenging problem due to both the low signal-to-noise ratio in financial markets (Gu et al.,, 2020) and the presence of volatility clustering. To this end, we present the Combined method for the simultaneous forecasting asset returns and modelling of forecast uncertainty in presence of volatility clustering. Our proposed method extends and simplifies the work of Lakshminarayanan et al., (2017) and Amini et al., (2020). We propose to use a SMD which uses a Gamma prior for scale $\nu$ uncertainty as a simpler alternative to the NIG prior which places a Normal prior to $\mu$ and an Inverse-Gamma prior to $\sigma^{2}$. Parameters of SMD are modelled using separate subnetworks. Together with model averaging and the use of log squared returns as inputs, we show that our proposed method can successfully model time-varying variance of the DGP, while providing superior return forecasting performance than the two state-of-the-art neural network UQ methods — Evidential and Ensemble. This is illustrated through the successful quantification of forecast uncertainty of two financial time-series datasets: cryptocurrency and U.S. equities.
@@
Our proposed SMD formulation offers an avenue to resolve some of the criticisms of Meinert et al., (2022) and Bengs et al., (2023). In particular, our SMD parameterization has three effective parameters and thus does not have any unresolved degrees of freedom. Setting $\alpha=\beta$ leads to a marginal t-distribution where the three distributional parameters ($\gamma,\sigma^{2},\alpha$) relate directly to the location, scale and shape of the t-distribution, without the need of a hierarchical model. In this formulation, epistemic uncertainty is assumed to be the difference between the predictive (t-distributed) and aleatoric (Normal-distributed) uncertainties. This assumption prioritizes a simpler model over granular attribution between aleatoric and epistemic uncertainties given by the NIG prior in Evidential. Moreover, Meinert et al., (2022) pointed out that the granular control comes at the cost of an unresolved degree of freedom. This also makes for a potential future research direction to evaluate such cost.
Despite assuming a more relaxed prior to estimate epistemic uncertainty, we show empirically that our method is able to accurately predict forecast errors, similar to the success that Evidential demonstrated in other real world applications (e.g., see Liu et al.,, 2021; Soleimany et al.,, 2021; Cai et al.,, 2021; Singh et al.,, 2022; Li and Liu,, 2022). From a finance application perspective, forecast uncertainty can be used to size bets, or as advanced warning to protect the portfolio from downside risk. For example, if forecast uncertainty reaches a certain threshold, an investor could purchase portfolio insurance (e.g., put options) or liquidate positions to reduce risk. The ability to attribute epistemic and aleatoric uncertainties may also allow for more advanced portfolio optimization techniques to be developed in future research (e.g., place different risk aversions on the two sources of uncertainties). Lastly, UQ in time-series applications is a relatively under-explored area of literature. We believe this paper can lead to further advancements of UQ in complex time-series. For example, the nonparametric quantile regression of Hüttel et al., (2023) can be applied to SMD by expressing the AL distribution hierarchically as scale mixtures of asymmetric uniforms with Gamma mixing density (Choy and Chan,, 2008; Wichitaksorn et al.,, 2015). Moreover, SMD also has a direct multivariate extension with conditional multivariate Normal and Gamma mixing densities to facilitate multivariate evidential regressions.
@@ Appendix A Marginal distribution of a Scale Mixture
From Equation (13), we have $y|\nu\sim\operatorname{N}(\gamma,\tfrac{\sigma^{2}}{\nu}),\ \nu\sim{\rm Gamma}(\alpha,\beta)$. Marginalizing over $\nu$ produces the data likelihood,
$\mathrm{p}(y|\gamma,\sigma^{2},\alpha,\beta)$ $=\int_{0}^{\infty}\mathrm{p}_{\operatorname{N}}(y|\gamma,\sigma^{2}\nu^{-1})\mathrm{p}_{\operatorname{G}}(\nu|\alpha,\beta)\mathop{}\!\mathrm{d}\nu$
$=\int_{0}^{\infty}\left[\sqrt{\frac{\nu}{2\pi\sigma^{2}}}\exp\left\{-\frac{\nu(y-\gamma)^{2}}{2\sigma^{2}}\right\}\right]\left[\frac{\beta^{\alpha}}{\Gamma(\alpha)}\nu^{\alpha-1}\exp^{-\beta\nu}\right]\mathop{}\!\mathrm{d}\nu$
$=\frac{\beta^{\alpha}}{\Gamma(\alpha)\sqrt{2\pi\sigma^{2}}}\int_{0}^{\infty}\nu^{\alpha-\tfrac{1}{2}}\exp\left\{-\frac{\nu(y-\gamma)^{2}}{2\sigma^{2}}-\beta\nu\right\}\mathop{}\!\mathrm{d}\nu$
$=\frac{\beta^{\alpha}}{\Gamma(\alpha)\sqrt{2\pi\sigma^{2}}}\left[\frac{(y-\gamma)^{2}}{2\sigma^{2}}+\beta\right]^{-(\alpha+\tfrac{1}{2})}\int_{0}^{\infty}\left\{\nu\left[\frac{(y-\gamma)^{2}}{2\sigma^{2}}+\beta\right]\right\}^{\alpha-\tfrac{1}{2}}$
$\quad\exp\left\{-\nu\left[\frac{(y-\gamma)^{2}}{2\sigma^{2}}+\beta\right]\right\}\mathop{}\!\mathrm{d}\left\{\nu\left[\frac{(y-\gamma)^{2}}{2\sigma^{2}}+\beta\right]\right\},$
@@
$=\frac{\Gamma(\frac{2\alpha+1}{2})}{\Gamma(\frac{2\alpha}{2})}\frac{1}{\sqrt{2\pi\sigma^{2}\beta}}\left[\frac{(y-\gamma)^{2}}{2\sigma^{2}\beta}+1\right]^{-(\frac{2\alpha+1}{2})}\hskip 119.50157pt$
@@
Appendix B Benchmarking on UCI dataset for non time-series
In this section, we compare Combined to Ensemble and Evidential using the UCI benchmark dataset. This is intended to facilitate a direct comparison to Lakshminarayanan et al., (2017) and Amini et al., (2020) for non-time-series UQ using the same dataset from both papers. This data set comes from a collection consisting of nine real world regression problems, each with 10–20 features and hundreds to tens of thousands of observations.
B.1 Comparison in UCI datasets
Table 4: Comparing Ensemble (Lakshminarayanan et al.,, 2017), Evidential (Amini et al.,, 2020) and Combined (this work) on RMSE and NLL using the UCI benchmark datasets. Results are averaged and standard deviations are calculated over 5 trials for each method. The best method for each dataset and metric are highlighted in bold.